# Rec-R1: Bridging Generative Large Language Models and User-Centric Recommendation Systems via Reinforcement Learning

### v2

@@ Abstract
We propose Rec-R1, a general reinforcement learning framework that bridges large language models (LLMs) with recommendation systems through closed-loop optimization. Unlike prompting and supervised fine-tuning (SFT), Rec-R1 directly optimizes LLM generation using feedback from a fixed, black-box recommendation model—without relying on synthetic SFT data from proprietary models like GPT-4o. This avoids the substantial cost and effort required for data distillation. To verify the effectiveness of Rec-R1, we evaluate Rec-R1 on two representative tasks: product search and sequential recommendation. Experimental results demonstrate that Rec-R1 not only consistently outperforms prompting- and SFT-based methods, but also achieves remarkable gains over strong discriminative baselines, even when used with simple retrievers like BM25. More impressively, Rec-R1 preserves the general-purpose capabilities of the LLM, in contrast to SFT, which often impairs instruction-following and reasoning. These findings suggest Rec-R1 as a promising foundation for continual task-specific adaptation without catastrophic forgetting.
@@ 1 Introduction
Recent advances in generative large language models (LLMs) have opened new possibilities for enhancing recommendation systems (Xi et al., 2024; Bao et al., 2023). Trained on massive web-scale corpora, LLMs possess extensive open-world knowledge, alongside strong capabilities in natural language understanding and reasoning (Achiam et al., 2023; Grattafiori et al., 2024; Yang et al., 2024; Lin et al., 2024). These strengths make LLMs particularly suitable for recommendation scenarios that involve language-driven inputs and implicit intent. As a result, recent studies have explored using LLMs in various stages of the recommendation pipeline, including query rewriting (Peng et al., 2024; Li et al., 2023b), user intent summarization (Torbati et al., 2023), and so on, to improving the performance of downstream recommendation tasks such as retrieval and ranking. These methods typically employ either zero- or few-shot prompting (Xi et al., 2024; Lyu et al., 2023; Ren et al., 2024; Li et al., 2023b) or supervised fine-tuning (SFT) (Luo et al., 2024b; Li et al., 2023a; Yang et al., 2023; Liao et al., 2024; Ji et al., 2024) to adapt LLMs to recommendation tasks.
However, most existing approaches still treat LLMs and recommendation models as disjoint components, with no closed feedback loop between LLM generation and recommendation performance (Peng et al., 2024; Zheng et al., 2023; Luo et al., 2024b; Li et al., 2023a; Yang et al., 2023; Liao et al., 2024). As a result, LLMs are typically optimized using proxy objectives rather than being directly trained using feedback from RecSys, which is often inconsistent with the ultimate goal of improving recommendation quality. Moreover, constructing high-quality supervision data for intermediate tasks—such as query rewriting—typically requires either human annotation, LLM APIs (e.g., GPT-4), or mining from historical interaction logs (Peng et al., 2024; Hou et al., 2024a). Nontheless, these sources rarely produce data that is truly aligned with optimal recommendation performance. Worse still, generating such data at scale is both time-consuming and financially expensive, especially when relying on human annotation or commercial LLMs. Figure 1 presents our proof-of-concept comparison that highlights the limitations of SFT in terms of both effectiveness and overhead.
@@
We evaluate Rec-R1 in two representative recommendation scenarios—product search and sequential recommendation—to demonstrate its effectiveness, though the framework itself is broadly applicable to a wider range of recommendation tasks (see Figure 4 and Appendix A). In product search, we observe that applying the Rec-R1 framework significantly improves overall recommendation performance, achieving state-of-the-art results on the evaluated benchmarks. In the sequential recommendation setting, Rec-R1 leads to consistent gains. More importantly, it shows strong performance in cold-start scenarios, where user profile information is absent, outperforming widely used sequential baselines. Beyond performance gains, we further investigate whether Rec-R1 preserves the general capabilities of the underlying language model. On the IFEval benchmark (Zhou et al., 2023), which measures instruction following capabilities, Rec-R1 maintains or even improves performance, while SFT causes a drop of over 27 points. This suggests that Rec-R1 enables task-specific adaptation without compromising the general-purpose capabilities of the initialized LLM. We summarize our main contributions as follows:
@@
  • We conduct extensive experiments on two representative recommendation tasks, i.e., product search and sequential recommendation, demonstrating that Rec-R1 significantly improves performance across different recommendation architectures. Notably, in product search, Rec-R1 improves the NDCG@100 score by up to 21.45 points for BM25-based retrievers, and by up to 18.76 points for dense discriminative models, compared to their respective baselines.
@@ 2.2 Theoretical Analysis of Existing Paradigms’ Limitations
This maximum likelihood estimation (MLE) objective encourages the learned policy $\pi_{\theta}$ to imitate $\pi_{g}$, but does not consider the downstream performance $f(a|s)$ during optimization. We now show that this training procedure imposes a fundamental performance ceiling, which we formalize in the following theorem.
Theorem 1 (SFT Converges Toward the Data-Generating Policy).
@@
[…]
@@
The proof can be found in Appendix C.1. Theorem 1 reveals a fundamental limitation of supervised fine-tuning: the learned policy $\pi_{\theta^{*}}$ is inherently constrained to imitate the data-generating policy $\pi_{g}$. Consequently, the recommendation performance of an SFT-trained model can at best approach—but never exceed—the performance of the policy used to generate the training data (e.g., GPT-4o). This theorem is empirically supported by our experiments on the ESCI dataset, as illustrated in Figure 1(a), where the performance of the SFT-trained model closely matches but does not surpass GPT-4o. However, as GPT-4o itself is not explicitly optimized for the downstream recommendation task, its performance is typically suboptimal.
@@ 2.3 The Rec-R1 Framework
Unlike SFT, this objective does not rely on manually labeled supervision or imitation of a fixed policy. Instead, it allows the model to continuously adapt its behavior to maximize performance on the downstream recommendation tasks. Following DeepSeek-R1 (Guo et al., 2025), we adopt Group Relative Policy Optimization (GRPO) Shao et al. (2024) to optimize the LLM policy. Compared to traditional algorithms such as PPO (Schulman et al., 2017), GRPO significantly reduces memory consumption during training while maintaining competitive performance. Further, we use rule-based reward functions derived from standard evaluation metrics (e.g., NDCG, Recall) rather than training a separate reward model, which helps mitigate reward hacking and avoids introducing additional biases.
Model Video Games Baby Products Office Products Sports and Outdoors
Sparse retrieval baselines
BM25 12.44 15.12 23.96 19.48
GPT-4o${}_{\text{+BM25}}$ 26.06 23.05 27.98 27.38
Qwen-2.5-3B-Instruct${}_{\text{+BM25}}$ 19.63 16.03 19.96 21.36
Dense retrieval baselines
[…]
GPT-4o+BLAIR-${}_{\text{+BLAIR-BASE}}$ 20.13 24.57 21.83 22.97
GPT-4o${}_{\text{+BLAIR-LARGE}}$ 23.99 24.10 22.99 24.67
Qwen-2.5-3B-Instruct+BLAIR-${}_{\text{+BLAIR-BASE}}$ 10.56 16.23 13.61 17.09
Qwen-2.5-3B-Instruct${}_{\text{+BLAIR-LARGE}}$ 17.34 16.10 17.29 17.74
Ours
[…]
(+21.45) (+14.15) (+10.65) (+12.44)
[…]
(+18.76) (+14.04) (+16.79) (+13.63)
[…]
(+15.53) (+12.80) (+12.95) (+14.19)
@@ 3 Experiments
To validate the effectiveness of Rec-R1, we conduct experiments on two representative recommendation scenarios: product search (§3.1) and sequential recommendation (§3.2). We also perform detailed analyses to examine the generalization ability after training (§3.3). More discussions and analysis can be found in Appendix F and case study in Appendix G.
@@ 3.1.2 Results
GPT-4o+BLAIR-${}_{\text{+BLAIR-BASE}}$ 21.12 19.54 17.22 22.68
GPT-4o${}_{\text{+BLAIR-LARGE}}$ 23.40 21.00 18.69 25.74
Qwen-2.5-3B-Instruct+BLAIR-${}_{\text{+BLAIR-BASE}}$ 15.82 18.07 14.34 16.96
Qwen-2.5-3B-Instruct${}_{\text{+BLAIR-LARGE}}$ 18.20 19.19 15.37 18.94
@@
Model Transductive Setting Inductive Setting
R@10 N@10 R@50 N@50 R@10 N@10 R@50 N@50
Text-aware SRec baselines
SASRec+BLAIR-${}_{\text{+BLAIR-BASE}}$ 3.72 1.55 7.81 2.44 0.20 0.90 1.40 0.35
SASRec${}_{\text{+BLAIR-LARGE}}$ 3.90 2.12 8.74 3.17 0.40 0.15 1.50 0.40
[…]
UniSRec+BLAIR-${}_{\text{+BLAIR-BASE}}$ 4.09 2.23 8.74 3.21 3.70 2.08 6.20 2.61
UniSRec${}_{\text{+BLAIR-LARGE}}$ 4.09 2.17 8.55 3.17 3.60 2.08 6.00 2.60
Query rewriting baselines
GPT-4o${}_{\text{+BM25}}$ 0.00 0.00 1.48 0.33 1.00 0.50 2.90 0.90
Qwen-2.5-3B-Instruct${}_{\text{+BM25}}$ 1.30 0.75 2.41 0.98 1.80 0.84 3.60 1.25
Ours
[…]
(+2.23) (+0.99) (+3.35) (+1.24) (+4.20) (+2.54) (+4.70) (+2.64)
@@ Appendix A Applying Rec-R1 to Diverse Recommendation Paradigms
Figure 4: Illustration of how generative LLMs are applied in recommender systems, following the taxonomy in Lin et al. (2025). The upper row shows the use of LLMs for feature engineering, including (1) Query Rewriting, where the LLM reformulates the input query to improve retrieval, and (2) User/Item-level Feature Augmentation, where the LLM encodes user or item information into richer textual representations as input to a downstream model. The lower row demonstrates the use of LLMs as Scoring/Ranking Functions, including (3) Closed-Set Item Generation, where the LLM ranks a given candidate list, and (4) Open-Set Item Generation, where the LLM directly generate candidate items and matches them to a product pool. Note that this figure primarily reflects the inference-time setting—thus all LLMs are frozen. Our proposed Rec-R1 is compatible with all paradigms shown here, and our experiments have already validated its effectiveness in three of them: Query Rewriting, Feature Augmentation, and Open-Set Item Generation.
Figure 4 provides a taxonomy of how generative LLMs are incorporated into recommender systems, adapted from Lin et al. (2025). The taxonomy is divided into two main categories: LLM for Feature Engineering and LLM as Scoring/Ranking Functions, each with two sub-paradigms. In our experiments, we validate the applicability of Rec-R1 across three of the four major paradigms:
  • Query Rewriting: In the product search task, we use BM25 as the retriever, and the LLM generates rewritten user queries to improve retrieval. This corresponds to the top-left panel of Figure 4 and demonstrates the effectiveness of Rec-R1 for textual query rewriting.
@@
Toward Closed-Set Item Generation.
Although our current experiments do not cover the Closed-Set Item Generation paradigm (bottom-left in Figure 4), Rec-R1 is naturally extendable to this setting. The LLM can be trained to rank or score candidates conditioned on user information through reward signals from ranking quality. Notably, prior work in this direction has largely relied on prompting (Zhang et al., 2023a; Hou et al., 2024b; Di Palma et al., 2023; Di Palma, 2023) or SFT (Yang et al., 2023; Luo et al., 2024b). However, as shown in our generalization experiments, such approaches may lead to catastrophic forgetting. In contrast, Rec-R1 maintains strong domain-specific performance without sacrificing general capabilities (§3.3), making it a more robust solution.
@@
Given its compatibility across paradigms and its ability to retain general-purpose capabilities, we believe Rec-R1 provides a strong foundation for continual, reinforcement-based alignment of LLMs with evolving recommendation goals. Future work will explore extending Rec-R1 into all four paradigms, ultimately enabling lifelong recommendation agents that adapt flexibly to new tasks and domains without retraining from scratch.

### v3

@@ Abstract
We propose Rec-R1, a general reinforcement learning framework that bridges large language models (LLMs) with recommendation systems through closed-loop optimization. Unlike prompting and supervised fine-tuning (SFT), Rec-R1 directly optimizes LLM generation using feedback from a fixed, black-box recommendation model—without relying on synthetic SFT data from proprietary models like GPT-4o. This avoids the substantial cost and effort required for data distillation. To verify the effectiveness of Rec-R1, we evaluate Rec-R1 on three representative tasks: product search, sequential recommendation, and product re-ranking. Experimental results demonstrate that Rec-R1 not only consistently outperforms prompting- and SFT-based methods, but also achieves remarkable gains over strong discriminative baselines, even when used with simple retrievers like BM25. More impressively, Rec-R1 preserves the general-purpose capabilities of the LLM, in contrast to SFT, which often impairs instruction-following and reasoning. These findings suggest Rec-R1 as a promising foundation for continual task-specific adaptation without catastrophic forgetting.
@@ 1 Introduction
Recent advances in generative large language models (LLMs) have opened new possibilities for enhancing recommendation systems (Xi et al., 2024; Bao et al., 2023). Trained on massive web-scale corpora, LLMs possess extensive open-world knowledge, alongside strong capabilities in natural language understanding and reasoning (Achiam et al., 2023; Grattafiori et al., 2024; Yang et al., 2024; Lin et al., 2024). These strengths make LLMs particularly suitable for user-centric recommendation scenarios, where the key challenge is to understand language-driven inputs, capture implicit user intent, and align recommendations with user needs. As a result, recent studies have explored using LLMs in various stages of the recommendation pipeline, including query rewriting (Peng et al., 2024; Li et al., 2023b), user intent summarization (Torbati et al., 2023), and so on, to improving the performance of downstream recommendation tasks such as retrieval and ranking. These methods typically employ either zero- or few-shot prompting (Xi et al., 2024; Lyu et al., 2023; Ren et al., 2024; Li et al., 2023b) or supervised fine-tuning (SFT) (Luo et al., 2024b; Li et al., 2023a; Yang et al., 2023; Liao et al., 2024; Ji et al., 2024) to adapt LLMs to recommendation tasks. Figure 1 illustrates the major paradigms where LLMs are applied in RecSys.
Figure 1: Illustration of how generative LLMs are applied in recommender systems (LLM4Rec), following the taxonomy in Lin et al. (2025). The upper row shows the use of LLMs for feature engineering, including (1) Query Rewriting, where the LLM reformulates the input query to improve retrieval, and (2) User/Item-level Feature Augmentation, where the LLM encodes user or item information into richer textual representations as input to a downstream model. The lower row demonstrates the use of LLMs as Scoring/Ranking Functions, including (3) Closed-Set Item Generation, where the LLM ranks a given candidate list, and (4) Open-Set Item Generation, where the LLM directly generate candidate items and matches them to a product pool. Note that this figure primarily reflects the inference-time setting—thus all LLMs are frozen. Our proposed Rec-R1 is compatible with all paradigms shown here (see Appendix A).
However, most existing approaches still treat LLMs and recommendation models as disjoint components, with no closed feedback loop between LLM generation and recommendation performance (Peng et al., 2024; Zheng et al., 2023; Luo et al., 2024b; Li et al., 2023a; Yang et al., 2023; Liao et al., 2024). As a result, LLMs are typically optimized using proxy objectives rather than being directly trained using feedback from RecSys, which is often inconsistent with the ultimate goal of improving recommendation quality. Moreover, constructing high-quality supervision data for intermediate tasks—such as query rewriting—typically requires either human annotation, LLM APIs (e.g., GPT-4), or mining from historical interaction logs (Peng et al., 2024; Hou et al., 2024a). Nonetheless, these sources rarely produce data that is truly aligned with optimal recommendation performance. Worse still, generating such data at scale is both time-consuming and financially expensive, especially when relying on human annotation or commercial LLMs. Figure 2 presents our proof-of-concept comparison that highlights the limitations of SFT in terms of both effectiveness and overhead.
@@
We evaluate Rec-R1 in three representative recommendation scenarios—product search, sequential recommendation, and product re-ranking—to demonstrate its effectiveness, though the framework itself is broadly applicable to a wider range of recommendation tasks (see Appendix A). In product search, we observe that applying the Rec-R1 framework significantly improves overall recommendation performance, achieving state-of-the-art results on the evaluated benchmarks. In the sequential recommendation setting, Rec-R1 leads to consistent gains. More importantly, it shows strong performance in cold-start scenarios, where user profile information is absent, outperforming widely used sequential baselines. In product re-ranking, Rec-R1 also shows clear superiority over both strong cross-encoder and LLM-based reranker baselines. Beyond performance gains, we further investigate whether Rec-R1 preserves the general capabilities of the underlying language model. On the IFEval benchmark (Zhou et al., 2023), which measures instruction following capabilities, Rec-R1 maintains or even improves performance, while SFT causes a drop of over 27 points. This suggests that Rec-R1 enables task-specific adaptation without compromising the general-purpose capabilities of the initialized LLM. We summarize our main contributions as follows:
@@
  • We conduct extensive experiments on three representative recommendation tasks, i.e., product search, sequential recommendation and product re-ranking, demonstrating that Rec-R1 significantly improves performance across different recommendation architectures. Notably, in product search, Rec-R1 improves the NDCG@100 score by up to 21.45 points for BM25-based retrievers, and by up to 18.76 points for dense discriminative models, compared to their respective baselines.
@@ 2.2 Theoretical Analysis of Existing Paradigms’ Limitations
This maximum likelihood estimation (MLE) objective encourages the learned policy $\pi_{\theta}$ to imitate $\pi_{g}$, but does not consider the downstream performance $f(a|s)$ during optimization. We now show that this training procedure imposes a fundamental performance ceiling, which we formalize in the following fact.
Fact 1 (SFT Converges Toward the Data-Generating Policy).
@@
    $\hat{p}(s,a)\xrightarrow{\text{a.s.}}p(s)\pi_{g}(a|s)\quad\text{as}\quad N\to\infty,$
@@
The proof can be found in Appendix C.1. Fact 1 reveals a fundamental limitation of supervised fine-tuning: the learned policy $\pi_{\theta^{*}}$ is inherently constrained to imitate the data-generating policy $\pi_{g}$. Consequently, the recommendation performance of an SFT-trained model can at best approach—but never exceed—the performance of the policy used to generate the training data (e.g., GPT-4o). This fact is empirically supported by our experiments on the ESCI dataset, as illustrated in Figure 2(a), where the performance of the SFT-trained model closely matches but does not surpass GPT-4o. However, as GPT-4o itself is not explicitly optimized for the downstream recommendation task, its performance is typically suboptimal.
@@ 2.3 The Rec-R1 Framework
Unlike SFT, this objective does not rely on manually labeled supervision or imitation of a fixed policy. Instead, it allows the model to continuously adapt its behavior to maximize performance on the downstream recommendation tasks. From the perspective of the optimization objective, Eq. 2 minimizes a proxy loss that indirectly relates to recommendation quality, whereas Eq. 5 directly optimizes the true recommendation objective, ensuring alignment with the downstream evaluation metrics. Following DeepSeek-R1 (Guo et al., 2025), we adopt Group Relative Policy Optimization (GRPO) (Shao et al., 2024) to optimize the LLM policy. Compared to traditional algorithms such as PPO (Schulman et al., 2017), GRPO significantly reduces memory consumption during training while maintaining competitive performance. Further, we use rule-based reward functions derived from standard evaluation metrics (e.g., NDCG, Recall) rather than training a separate reward model, which helps mitigate reward hacking and avoids introducing additional biases.
@@ 3 Experiments
To validate the effectiveness of Rec-R1, we conduct experiments on three representative recommendation scenarios: product search (§3.1), sequential recommendation (§3.2), and product re-ranking (§3.3). We also perform detailed analyses to examine the generalization ability after training (§3.4). More discussions and analysis can be found in Appendix F and case study in Appendix G.
@@ 3.1.1 Experimental Setup
Ours
$\text{{Rec-R1}-3B}_{\text{+BM25}}$ 33.89 29.27 34.61 31.92
(+21.45) (+14.15) (+10.65) (+12.44)
$\text{{Rec-R1}-3B}_{\text{+BLAIR-BASE}}$ 28.51 29.24 33.98 30.71
(+18.76) (+14.04) (+16.79) (+13.63)
$\text{{Rec-R1}-3B}_{\text{+BLAIR-LARGE}}$ 31.41 28.76 34.12 32.49
(+15.53) (+12.80) (+12.95) (+14.19)
3.1.2 Results
Model Video Games Baby Products Office Products Sports and Outdoors
Sparse retrieval baselines
BM25 7.82 6.39 8.70 7.28
@@ 3.1.2 Results
[…]
Dense retrieval baselines
[…]
[2pt/2pt]
[…]
[2pt/2pt]
[…]
@@ 3.2.1 Experimental Setup
Model Transductive Setting Inductive Setting
R@10 N@10 R@50 N@50 R@10 N@10 R@50 N@50
Text-aware SRec baselines
[…]
[2pt/2pt]
[…]
Query rewriting baselines
[…]
Ours
$\text{{Rec-R1}-3B}_{\text{+BM25}}$ 3.53 1.74 5.76 2.22 6.00 3.38 8.30 3.89
(+2.23) (+0.99) (+3.35) (+1.24) (+4.20) (+2.54) (+4.70) (+2.64)
@@ 3.2.2 Results
3.3 Product Re-ranking
3.3.1 Experimental Setup
Task Definition. In the product re-ranking task, the input is a natural language query and an initial candidate list of $m$ items, denoted as $s=(q,C)\in\mathcal{S}$ where $q$ is the user query and $C=[i_{1},\dots,i_{m}]$ is the candidate set. The goal is to reorder $C$ so that items more relevant to $q$ are ranked higher. Given $s$, the LLM produces $a\in\mathcal{A}$, a re-ranked list of the $m$ candidates (i.e., a permutation of $C$), denoted by $a=[i_{\pi(1)},\dots,i_{\pi(m)}]\in\mathcal{A}$, where $\pi$ is a permutation function over $\{1,\dots,m\}$. To evaluate performance, we define a relevance dictionary $\mathcal{D}$ that maps each input $s$ to corresponding ground-truth item list. The reward score $f(a|s)\in\mathbb{R}$ is computed by comparing the re-ranked list $a$ against the target item list $\mathcal{D}(s)$. We use NDCG@10 as the ranking metric.
Dataset and Baselines. We conduct experiments on the ESCI dataset (Reddy et al., 2022). The candidate items $C$ are retrieved using the Rec-R1-Retriever given the user query, and subsequently used for the re-ranking stage. To ensure fair comparison, all re-ranking models operate on the same candidate lists obtained from the Rec-R1-Retriever. We compare Rec-R1 with two families of baselines: (1) traditional cross-encoder re-rankers (Zhang et al., 2025), where we adopt the state-of-the-art open-source Qwen3-Reranker family, including the 4B and 8B variants; and (2) LLM-based re-rankers, including GPT-4o and Qwen2.5-3B-Instruct, which directly generate ranking orders conditioned on the query and candidate list. Further implementation details and hyperparameters are provided in Appendix E.3.
3.3.2 Results
Table 4: Performance comparison of different re-rankers on ESCI. We report the NDCG@10 scores across four domains. The best performance is denoted in bold. The numbers in gray indicate the improvement of Rec-R1 over its corresponding initialized model, i.e., Qwen2.5-3B-Instruct.
Model Video Games Baby Products Office Products Sports and Outdoors
Cross-encoder re-rankers
Qwen3-Reranker-4B 44.29 46.41 48.06 47.42
Qwen3-Reranker-8B 44.28 44.95 47.73 44.51
LLM-based re-ranker baselines
Qwen2.5-3B-Instruct 5.38 9.21 4.26 4.37
GPT-4o 61.57 56.18 53.62 53.62
Ours
Rec-R1-3B 53.04 57.08 60.91 62.91
(+47.66) (+47.87) (+56.65) (+58.54)
Table 4 shows the re-ranking performance of different models on the ESCI dataset. From the table, Rec-R1 consistently achieves the best overall performance across all domains. Notably, compared with the initialized Qwen2.5-3B-Instruct model, Rec-R1 delivers very large improvements. Furthermore, against SoTA open-source embedding/reranker models (Qwen3-Reranker family), Rec-R1 demonstrates substantial superiority. These results confirm that Rec-R1 ’s optimization framework can also be applied to the closed-set item generation paradigm within LLM4Rec, further supporting our claim that Rec-R1 provides a general way to bridge LLMs and RecSys.
@@ Appendix A Applying Rec-R1 to Diverse Recommendation Paradigms
Figure 1 provides a taxonomy of how generative LLMs are incorporated into recommender systems, adapted from Lin et al. (2025). The taxonomy is divided into two main categories: LLM for Feature Engineering and LLM as Scoring/Ranking Functions, each with two sub-paradigms. In our experiments, we validate the applicability of Rec-R1 across all of the four major paradigms:
  • Query Rewriting: In the product search task, we use BM25 as the retriever, and the LLM generates rewritten user queries to improve retrieval. This corresponds to the top-left panel of Figure 1 and demonstrates the of Rec-R1 for textual query rewriting.
@@
  • Closed-Set Item Generation: In the product re-ranking task, the LLM is trained to rank or score candidates conditioned on user query and initialized product candidates through reward signals from ranking quality. This aligns with the bottom-left panel in Figure 1. Notably, prior work in this direction has largely relied on prompting (Zhang et al., 2023a; Hou et al., 2024b; Di Palma et al., 2023; Di Palma, 2023) or SFT (Yang et al., 2023; Luo et al., 2024b). However, as shown in our generalization experiments, such approaches may lead to catastrophic forgetting. In contrast, Rec-R1 maintains strong domain-specific performance without sacrificing general capabilities (§3.4), making it a more robust solution.
@@
Given its compatibility across paradigms and its ability to retain general-purpose capabilities, we believe Rec-R1 provides a strong foundation for continual, reinforcement-based alignment of LLMs with evolving recommendation goals. Future work will explore extending Rec-R1 into more tasks, ultimately enabling lifelong recommendation agents that adapt flexibly to new tasks and domains without retraining from scratch.